OpenStudy (anonymous):

The length of a rectangle is 6 more than 2 times a number. The width is the number plus 3 and the area is 32. What is the width of the rectangle?

OpenStudy (anonymous):

L = 2n+6 W = n+3 A = 32 = L*W can you do the rest?